# SDE-based Monte Carlo dose calculation for proton therapy validated against Geant4

### v2

@@ 1 Introduction
Rather than relying exclusively on increasing computational power, recent efforts have also focused on advancing the underlying physical and mathematical models to improve efficiency at the algorithmic level. In addition to variance reduction techniques such as the Leq formalism and related approaches (1), which accelerate Monte Carlo simulations by reducing statistical noise, there is growing interest in alternative formulations of the transport problem itself.
@@ 2.2.2 Test cases
For all cases, a monoenergetic proton beam (energy of 100 or 150 MeV, nozzle radius of 5 mm, energy spread $\sigma_{E}$ of 0.1 keV and radial dispersion $\sigma_{r}$ of 0.05 rad) was directed perpendicularly into water, with no air gap between the source and the phantom. The chemical compositions and densities for all materials are the same for Geant4 and the SDE model. For all cases, $1\times 10^{6}$ protons were fired per run, and dose was scored in both cases using a 3D grid of 1 mm3 voxels across the irradiated volume, enabling pointwise comparison of the resulting dose distributions. The SDE model requires angular limits for backscatter and Rutherford events, which were set at 0.04 and 0.02 rad, respectively. Since the SDE model employs a fixed step length, this was set to 0.5 mm for all simulations. For consistency, the maximum step length in Geant4 was constrained to the same value to ensure a fair comparison in terms of time performance. For both simulators, the energy deposition is assigned at the step midpoint to minimise bias in regions of steep gradients.
@@ 3.2 Longitudinally heterogeneous slab phantom
Figure 6: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a water phantom with a 1-cm bone layer and a 2-cm low density lung layer. (a) Integral depth-dose curves, (b) Central-axis depth-dose curves. Lower subplots show pointwise percentage dose differences.
@@ 3.3 Laterally heterogeneous composite phantom
Having established good agreement between the SDE model and Geant4 for the homogeneous and slab phantom configurations, the laterally heterogeneous insert phantom provides a more stringent test of the model’s response to localised lateral heterogeneities, which are known to challenge reduced proton transport models (Schaffner et al., 1999). Figure 9 shows the integrated and central-axis dose distributions, which remain in close agreement and within the same percentage difference ranges observed for the previous test cases.
Figure 9: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a heterogeneous phantom with a 2-cm off-axis bone insert. (a) Integral depth-dose curves, (b) Central-axis depth-dose curves. Lower subplots show pointwise percentage dose differences.
Both models reproduce the two expected high-dose peaks arising from the range shift experienced by the upper half of the beam. The agreement in proton range for the more prominent peak is within 0.2 mm for the 100 MeV case and within 0.5 mm for the 150 MeV case, with a comparable level of agreement retained for the secondary peak.
@@
Gamma analysis for the insert phantom yielded pass rates of 97% for the 100 MeV beam and 98.5% for the 150 MeV beam, using the same stringent criteria applied in the previous scenarios. In terms of computational performance, the SDE model remains between 2.5 and 2.9 times faster than single-threaded Geant4 for this configuration.

### v3

@@ 1 Introduction
Rather than relying exclusively on increasing computational power, recent efforts have also focused on advancing the underlying physical and mathematical models to improve efficiency at the algorithmic level. In addition to variance reduction techniques such as the Leq formalism and related approaches (Maneval et al., 2018), which accelerate Monte Carlo simulations by reducing statistical noise, there is growing interest in alternative formulations of the transport problem itself.
@@ 2.2.2 Test cases
In the heterogeneous phantom simulations, the bone and lung regions were defined using standard material definitions. The bone insert was modelled using the Geant4 material G4_B-100_BONE with a density of 1.45 g/cm3, while the lung region was modelled using the ICRP 110 reference human phantom composition with a density of 0.385 g/cm3. These material definitions were used consistently in both the Geant4 reference simulations and the SDE model parameterisation.
For all cases, a monoenergetic proton beam (energy of 100 or 150 MeV, nozzle radius of 5 mm, energy spread $\sigma_{E}$ of 0.1 keV and radial dispersion $\sigma_{r}$ of 0.05 rad) was directed perpendicularly into water, with no air gap between the source and the phantom. The chemical compositions and densities for all materials are the same for Geant4 and the SDE model. For all cases, $1\times 10^{6}$ protons were fired per run, and dose was scored in both cases using a 3D grid of 1 mm3 voxels across the irradiated volume, enabling pointwise comparison of the resulting dose distributions. The dose in each voxel was calculated as the deposited energy divided by the local voxel mass, corresponding to dose-to-medium. The SDE model requires angular limits for backscatter and Rutherford events, which were set at 0.04 and 0.02 rad, respectively. Since the SDE model employs a fixed step length, this was set to 0.5 mm for all simulations. For consistency, the maximum step length in Geant4 was constrained to the same value to ensure a fair comparison in terms of time performance. For both simulators, the energy deposition is assigned at the step midpoint to minimise bias in regions of steep gradients.
@@ 3.2 Longitudinally heterogeneous slab phantom
Figure 6: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a water phantom with a 1-cm bone layer and a 2-cm low density lung layer, with material boundaries delimited with grey solid lines. (a) Integral depth-dose curves, (b) Central-axis depth-dose curves. Lower subplots show pointwise percentage dose differences.
@@ 3.3 Laterally heterogeneous composite phantom
Having established good agreement between the SDE model and Geant4 for the homogeneous and slab phantom configurations, the laterally heterogeneous insert phantom provides a more stringent test of the model’s response to localised lateral heterogeneities, which are known to challenge reduced proton transport models (Schaffner et al., 1999). Figure 9 shows the integrated and central-axis dose distributions, which remain in close agreement and within the same percentage difference ranges observed for the previous test cases. Both models reproduce the two expected high-dose peaks arising from the range shift experienced by the upper half of the beam. The agreement in proton range for the more prominent peak is within 0.2 mm for the 100 MeV case and within 0.5 mm for the 150 MeV case, with a comparable level of agreement retained for the secondary peak.
Figure 9: 1D comparison between SDE and Geant4 for two monoenergetic proton beams (100 and 150 MeV) in a laterally heterogeneous phantom with a 2-cm off-axis bone insert, with material boundaries delimited with grey solid lines. (a) Integral depth-dose curves, (b) Central-axis depth-dose curves. Lower subplots show pointwise percentage dose differences.
@@
Gamma analysis for the insert phantom yielded pass rates of 96.2% for the 100 MeV beam and 98.2% for the 150 MeV beam, using the same stringent criteria applied in the previous scenarios. In terms of computational performance, the SDE model remains between 2.5 and 2.9 times faster than single-threaded Geant4 for this configuration.
@@ 4 Discussion and Conclusion
Additionally, it is worth mentioning that the reported dose corresponds to dose-to-medium, obtained by dividing the deposited energy by the local voxel mass. In heterogeneous materials, dose-to-medium differs from dose-to-water due to material-dependent stopping powers and nuclear interactions, as discussed by Paganetti (2009). In the present study, all comparisons between the SDE model and Geant4 are performed consistently using dose-to-medium. For potential use within treatment planning workflows, conversion to dose-to-water could be performed using established stopping-power–based formalisms such as that of Paganetti.